OpenStudy (anonymous):

FInd the zeros of the function using Fundamental Theorem of Algebra, the Rational Root Theorem, Descartes' Rule of Signs, and the Factor Theorem. x^2 - 7x - 72 Please try your best to explain how you do this or just show your work, it will help me with the rest of the problems.

OpenStudy (anonymous):

that function has not rational zeros

OpenStudy (anonymous):

That's what I came up with too, what about complex zeros though?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

the zeros are real

OpenStudy (anonymous):

the problem says there can be complex zeros, i just have trouble finding them

OpenStudy (anonymous):

the zeros r real just use a calculator or the quadratic formual and u will see

OpenStudy (anonymous):

What i had to do for this assignment, is take a rectangular box, measure the lwh and find the volume: l:7 w:4 h:3 V:84 Then I had to rewrite the equation as if I were solving for length, with the volume plugged in: 84 = (4)(3)x

OpenStudy (anonymous):

84 = 12x Then I had to rewrite the equation expressing the width and height of the object in terms of x plus or minus a constant. For example, if the height measurement is 4 inches longer than the length, then the expression for the height will be (x + 4), and I got: 84 = (x - 3)(x - 4) Then I had to write in standard form: 84 = x^2 - 4x - 3x +12 84 = x^2 - 7x + 12 0 = x^2 - 7x - 72 And then I had to write it as a function (f(x)): f(x) = x^2 - 7x - 72 Now I'm stuck trying to finish this, by finding the zeros of the function!

OpenStudy (rajee_sam):

You missed out an x there. The equation should have been : 84 = (x-3)(x-4) x . not just (x-3)(x-4) . correct that and try again.

OpenStudy (anonymous):

Oh wow! Thanks for that, I'll try again! And sorry about the mistake both of you!

OpenStudy (rajee_sam):

You started with a real world object so you should have real numbers only as your answer.

OpenStudy (anonymous):

okay, thanks!

OpenStudy (anonymous):

I should get this once that was simplified? 0 = x^3 - 7x^2 + 12 - 84